40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10615     Entry time: Thu Oct 16 03:13:23 2014
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRMI on REFL165, and more 

 The first thing I looked at tonight was locking the PRMI on REFL 165.

I locked the PRMI (no arms), and checked the REFL 165 demod phase. I also found the input matrix configuration that allowed me to acquire PRMI lock directly on REFL165.  After locking the arms on ALS, I tried to lock the PRMI with REFL 165 and failed.  So, I rechecked the demod phase and the relative transfer functions between REFL 165 and REFL 33.  The end of the story is that, even with the re-tuned demod phase for CARM offset of a few nanometers, I cannot acquire PRMI lock on REFL 165, nor can I transition from REFL 33 to REFL 165.  We need to revisit this tomorrow.

IFO configuration CARM offset [cts] REFL 165 demod phase [deg]
Found as-is N/A +145
PRMI, no arms N/A -135
PRFPMI +3 +110
PRFPMI +2 +110
PRFPMI +1 +110
PRFPMI +0.5 +120

 

IFO configuration REFL 33 I / REFL 165 I (PRCL) REFL 33 Q / REFL 165 Q (MICH)
PRMI, no arms +0.1 +0.22, although easier to acquire lock with +0.1
PRFPMI, CARM offset = +3 -0.09  (TF measured, no lock) +0.033  (TF measured, no lock)

For the PRMI-only case, I ended up using 0.1's in the input matrix, and I added an FM 1 to the MICH filter bank that is a flat gain of 2.2, and then I had it trigger along with FM2.

I turned this FM1 off (and no triggering) while trying to transition from REFL33 to REFL165 in the PRFPMI case, but that didn't help.  I think that maybe I need to remeasure my transfer functions or something, because I could put values into the REFL165 columns of the input matrix while REFL33 was still 1's, but I couldn't remove (even if done slowly) the REFL33 matrix elements without losing lock of the PRMI.  So, we need to get the input matrix elements correct.


I also recorded some time series for a quick RAM investigation that I will work on tomorrow.  

I left the PRM aligned, but significantly misaligned both ITMs to get data at the REFL port of the RAM that we see.  I also aligned the PRMI (no arms) and let it flash so that I can see the pk-pk size of our PDH signals.  I need to remember to calibrate these from counts to meters.  

Raw data is in /users/jenne/RAM/ .


I have not tried any new DARM signals, since PRMI wasn't working with 3f2.  

We should get to that as soon as we fix the PRMI-3f2 situation.

ELOG V3.1.3-